OpenStudy (anonymous):

\[f(x)=4(x+1)^{2}-2\] x-intercept of parabola?

myininaya (myininaya):

aimstre just answered this lol

OpenStudy (anonymous):

\[(4)(x+1)^2-2=0\] \[4(x+1)^2=2\] \[(x+1)^2=\frac{1}{4}\] \[x+1=\pm\sqrt{\frac{1}{4}}=\pm\frac{1}{2}\]

OpenStudy (anonymous):

\[x=-1+\frac{1}{2}=-\frac{1}{2}\]

myininaya (myininaya):

oh no 2/4=1/2 not 1/4

OpenStudy (anonymous):

or \[x=-1-\frac{1}{2}= -\frac{3}{2}\]

OpenStudy (anonymous):

?

myininaya (myininaya):

(x+1)^2=1/2

OpenStudy (anonymous):

ooooooooooooooooops

OpenStudy (anonymous):

sorry myininaya is right and i am wrong

myininaya (myininaya):

:) i like to hear that more

OpenStudy (anonymous):

\[(x+1)^2=\frac{1}{2}\]

OpenStudy (anonymous):

this is not right...

OpenStudy (anonymous):

\[x+1=\pm\sqrt{\frac{1}{2}}\]

OpenStudy (anonymous):

\[x=-1\pm\frac{1}{\sqrt{2}}\]

myininaya (myininaya):

gj satellitle i knew you could do it

OpenStudy (amistre64):

yay!! lol

OpenStudy (anonymous):

hope so! i just wrote that 2/4 = 1/4 instead of1/2!

myininaya (myininaya):

lol

OpenStudy (anonymous):

must be time for some holy wine!

myininaya (myininaya):

maybe it does and we just have been doing math wrong all these years

myininaya (myininaya):

satellite the bat signal is out i mean the satellite signal is out

OpenStudy (anonymous):

could be. btw to tinka, there are many ways to write \[1+\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

or even \[-1+\frac{1}{\sqrt{2}}\]

myininaya (myininaya):

someone is requesting you

OpenStudy (anonymous):

me?

myininaya (myininaya):

top post